# A splitting method for numerical relativistic magnetohydrodynamics

### v1

@@ 3.3.1 Modified 2nd order TVD weights
Any linear combination of these interpolants ensures 2nd order spatial accuracy in smooth regions of numerical solution. Falle (1991) used a TVD slope limiter which is equivalent [^2] Falle (1991) also use the polynomial $P_{0}(x)=u_{i}$, for the case where $\beta_{+}\beta_{-}\leq 0$. to using following linear combination of the polynomials $P_{\pm}$
@@
3.3.2 3rd-order WENO weights
3rd-order WENO interpolation utilises the fact that the linear interpolation (46) yields the same value at $x=x_{i+1/2}$ as the quadratic interpolation based on the all three points of the stencil $S$ if $w_{-}=1/4$ and $w_{+}=3/4$, and the same value at $x=x_{i-1/2}$ if $w_{-}=3/4$ and $w_{+}=1/4$. Thus, two linear interpolants of the form (46), one per each interface of the cell, can be used to achieve 3rd-order accurate interpolation to the both interfaces. $\gamma_{a}=1/4$ and $\gamma_{b}=3/4$ are known as the ideal or linear weights. We denote the interpolant used for the interpolation to the $x_{i-1/2}$ interface of $i$-th cell as
@@
One may put $w_{+}^{r}=w_{-}^{l}=\gamma_{b}$ and $w_{+}^{l}=w_{-}^{r}=\gamma_{a}$, but this will lead to violent oscillations at shocks. Instead, WENO weights are nonlinear, reducing to the ideal weights only on very smooth solutions. At shocks, the linear interpolant with lower gradient should dominate. Since the 2nd-order TVD interpolation, described earlier, is also based on the three-point stencil $S$, has exactly the same form as the 3rd-order WENO interpolants, and already has the required behaviour at shocks, a mapping of the TVD weights, which is closed to the identity mapping at shocks but yields ideal weights on smooth solutions, suggests itself.
@@
$w_{+}^{l}\to\gamma_{a},\quad w_{+}^{r}\to\gamma_{b}\quad\mbox{as}\quad w_{+}\to 0.5\,,$ (57)
@@
$w_{+}^{r}=\gamma_{b}\alpha(w_{+})\,,$ (59)
@@
$\alpha(0)=0\,,\quad\alpha(0.5)=1\,,\quad\alpha(1)=1/\gamma_{b}\,.$ (60)
@@
To determine $\alpha(x)$ for $x\in[0.5,1]$, we require the function $w_{-}^{r}(w_{-})$ to be the same as $w_{+}^{r}(w_{+})$, apart from $\gamma_{b}$ replaced by $\gamma_{a}$, and write
$w_{-}^{r}=\gamma_{a}\alpha(w_{-})\,.$ (62)
@@
$w_{+}^{r}=1-\gamma_{a}\alpha(1-w_{+})\,.$
@@
$w_{+}^{r}(w_{+})$ =$=\begin{cases}\gamma_{b}p_{n}(w_{+}),&0\leq w_{+}\leq 0.5\,,\\ 1-\gamma_{a}p_{n}(1-w_{+}),&0.5<w_{+}\leq 1\,,\end{cases}$ (63)
@@
$w_{+}^{l}(w_{+})$ =$=\begin{cases}\gamma_{a}p_{n}(w_{+}),&0\leq w_{+}\leq 0.5\,,\\ 1-\gamma_{b}p_{n}(1-w_{+}),&0.5<w_{+}\leq 1\,,\end{cases}$ (65)
@@
3.3.3 Downgrading to 2rd-order TVD interpolation at strong shocks

### v3

@@ A.1 Modified 2nd order TVD weights
Any linear combination of these interpolants ensures 2nd order spatial accuracy in smooth regions of numerical solution. Falle (1991) used a TVD slope limiter which is equivalent [^2] Falle (1991) also use the polynomial $P_{0}(x)=u_{i}$, for the case where $\beta_{+}\beta_{-}\leq 0$. using following linear combination of the polynomials $P_{\pm}$
@@
A.2 3rd-order WENO weights
3rd-order WENO interpolation utilises the fact that the linear interpolation (116) yields the same value at $x=x_{i+1/2}$ as the quadratic interpolation based on the all three points of the stencil $S$ if $w_{-}=1/4$ and $w_{+}=3/4$, and the same value at $x=x_{i-1/2}$ if $w_{-}=3/4$ and $w_{+}=1/4$. Thus, two linear interpolants of the form (116), one per each interface of the cell, can be used to achieve 3rd-order accurate interpolation to the both interfaces. $\gamma_{\mbox{\tiny a}}=1/4$ and $\gamma_{\mbox{\tiny b}}=3/4$ are known as the ideal or linear weights. We denote the interpolant used for the interpolation to the $x_{i-1/2}$ interface of $i$-th cell as
@@
One may put $w_{+}^{r}=w_{-}^{l}=\gamma_{\mbox{\tiny b}}$ and $w_{+}^{l}=w_{-}^{r}=\gamma_{\mbox{\tiny a}}$, but this will lead to violent oscillations at shocks. Instead, WENO weights are nonlinear, reducing to the ideal weights only on very smooth solutions. At shocks, the linear interpolant with lower gradient should dominate. Since the 2nd-order TVD interpolation, described earlier, is also based on the three-point stencil $S$, has exactly the same form as the 3rd-order WENO interpolants, and already has the required behaviour at shocks, a mapping of the TVD weights, which is closed to the identity mapping at shocks but yields ideal weights on smooth solutions, suggests itself.
@@
$w_{+}^{l}\to\gamma_{\mbox{\tiny a}},\quad w_{+}^{r}\to\gamma_{\mbox{\tiny b}}\quad\mbox{as}\quad w_{+}\to 0.5\,,$ (127)
@@
$w_{+}^{r}=\gamma_{\mbox{\tiny b}}\alpha(w_{+})\,,$ (129)
@@
$\alpha(0)=0\,,\quad\alpha(0.5)=1\,,\quad\alpha(1)=1/\gamma_{\mbox{\tiny b}}\,.$ (130)
@@
To determine $\alpha(x)$ for $x\in[0.5,1]$, we require the function $w_{-}^{r}(w_{-})$ to be the same as $w_{+}^{r}(w_{+})$, apart from $\gamma_{\mbox{\tiny b}}$ replaced by $\gamma_{\mbox{\tiny a}}$, and write
$w_{-}^{r}=\gamma_{\mbox{\tiny a}}\alpha(w_{-})\,.$ (132)
@@
$w_{+}^{r}=1-\gamma_{\mbox{\tiny a}}\alpha(1-w_{+})\,.$
@@
$w_{+}^{r}(w_{+})$ =$=\begin{cases}\gamma_{\mbox{\tiny b}}p_{n}(w_{+}),&0\leq w_{+}\leq 0.5\,,\\ 1-\gamma_{\mbox{\tiny a}}p_{n}(1-w_{+}),&0.5<w_{+}\leq 1\,,\end{cases}$ (133)
@@
$w_{+}^{l}(w_{+})$ =$=\begin{cases}\gamma_{\mbox{\tiny a}}p_{n}(w_{+}),&0\leq w_{+}\leq 0.5\,,\\ 1-\gamma_{\mbox{\tiny b}}p_{n}(1-w_{+}),&0.5<w_{+}\leq 1\,,\end{cases}$ (135)
@@
A.3 Downgrading to 2rd-order TVD interpolation at strong shocks